# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v2

@@ 5.4. Multi-Precision INT-PE Array
Accumulation Stage. It receives the MUL-stage output (Res) and also iAcc from either ReCoN or the previous PE row. The controller signals the ADD stage via Outlier_Present to modify accumulation behavior based on whether inlier/outlier weights are present in the PE. If the Res corresponds to a multiplication of inlier weight with iAct, then the corresponding Res is added with iAcc using two $2$-bit $+$ $4$-bit adders with a multiplexer that enables multi-precision by propagating the carry from one adder to other. In low precision MODE, the two adders work independently to calculate partial sums in parallel. For high precision MODE, both adders work together to produce a single accumulation result with carry propagation through the multiplexer. On the other hand, if Res corresponds to multiplication of either of the outlier halves with iAct, the actual outlier accumulation is offloaded to ReCoN by concatenating Res and iAcc666the concatenated output is notated as $O_{\texttt{Upper},\texttt{Res}}$ or $O_{\texttt{Lower},\texttt{Res}}$.. This is done to prevent incorrect outlier partial sum calculation. The controller finally directs the partial sum output to ReCoN or the next PE row, via the OAcc_NoC/PE signal.
Figure 7. End-to-end example illustrating the working of MicroScopiQ with a $4\times 4$ PE array (only two rows shown)
@@ 5.5. Redistribution and Coordination NoC
The Redistribution and Coordination NoC (ReCoN) is a multistage butterfly NoC in the MicroScopiQ accelerator time-multiplexed and shared across all PE rows (see section 6.5). This is a more cost effective way of handling outliers compared to OliVe and GOBO (Guo et al., 2023; Zadeh et al., 2020) that handle outliers within the PE. Since PEs are large in number OliVe and GOBO (Guo et al., 2023; Zadeh et al., 2020) will incur larger costs compared to MicroScopiQ. Note: Not all rows require access to ReCoN, only rows that have one or more outlier weights in the PE require access. Since outliers are still minimal (though important) compared to inliers, time-multiplexing access to ReCoN is a cost-effective solution with minimal access conflicts to ReCoN (see section 6.5).
@@
ReCoN Input Interface. ReCoN is shared and time-multiplexed across all PE rows, and at a particular instant many rows might require access to it. At the inputs, the column-wise multiplexers guided by control signals from the controller, coordinate and resolve contention between different PE rows and ensure a fair scheduling. Due to the skewed data flow of systolic array computations (Jouppi et al., 2017; Ramachandran et al., 2024b), partial sums from different columns of the same row do not arrive at the ReCoN input at the same cycle. A synchronization buffer is positioned between ReCoN and column-wise multiplexers which propagates the partial sums through sets of buffers of different lengths (column producing the partial sum fastest has the most buffers), synchronizing the arrival of inputs from different columns. The synchronization buffer sends an ACK signal to the PE row whose values have been accepted so that it can progress in its computation and the other simultaneously contesting rows will be on hold and acknowledged in the next cycle. When there are N simultaneously contesting rows, the row that is acknowledged the last adds an N-1 cycle latency to its processing. As we demonstrate in Figure 11(b) and Figure 14(a) the number of row conflicts in accessing ReCoN are minimal and if latency is a critical factor multiple ReCoN units can be used to minimize conflicts.
ReCoN switch. It is a {2-input, 2-output} switch capable of handling multi-precision and performs three major operations based on a 3-bit configuration provided by the controller:
@@ 5.6. Post Processing and Output Scale Factor Computation
The shared output scale factor for $B_{M}$=128 is calculated as ${oAct_{sf}}={O_{sf}+iAct_{sf}}$ (See expansion of $O_{sf}$ above.). The calculation of $oAct_{sf}$ (with simple adders and subtractors) is independent of the processing done in the PE array. Therefore, we overlap the computation of $oAct_{sf}$ with the processing of oActs to efficiently hide the computation latency similar to prior works (Dai et al., 2021; Ramachandran et al., 2024b). Since we maintain different scales for inliers and outliers (see section 4.2), and the oAct scale factor is calculated based on the outlier scale factor ($O_{sf}$) (see above) (we found this to result in least quantization error of oActs), the oActs which are generated through computation with only inlier weights (identified through the in/out control signal) are shifted by the shift value depicted in the “Post Processing and Scale Compute Unit” in Figure 5 to ensure conformity with the final scale factor. Finally, the oActs are scaled with the computed output scale factor through a simple right shift (since it is a power-of-two scale factor, division can be implemented through right shift operation) and quantized to MX-INT-(4/8)128 before being sent to external memory or routed back to the iAct buffer for computation with the next layer’s weights. The cyclic behavior diagram of the post processing unit in Figure 5 depicts the single clock cycle latency incurred to scale, quantize and store the oActs. This is inline with SoTA accelerators’(Keller et al., 2023; Dai et al., 2021; Shao et al., 2019) post processing latency. Similar to prior works (Ramachandran et al., 2024b), (Keller et al., 2023) the post processing unit is also responsible for handling all non-linear operations.
@@
MicroScopiQ (Ours) 4/4 11.72 9.75 11.47 8.29 5.48 9.53 5.23 6.01 7.54 6.37
[…]
@@
MicroScopiQ (Ours) 2/8 11.94 9.98 9.06 8.06 6.59 9.17 6.53 6.27 7.98 7.02
[…]
Table 3. LLM quantization performance comparison of MicroScopiQ enhanced with OmniQuant (OQ).
@@
MicroScopiQ + OQ 2/8 7.12 5.74 7.21
Table 4. Comparison of quantization performance for LLaMA2-70B on various LLM benchmarks.
@@
MicroScopiQ 2/16 51.30 80.60 62.70 77.80
5.7. MicroScopiQ in Action
In this section, we discuss an end-to-end example that concretely showcases the working of MicroScopiQ. In the example shown in Figure 7, we employ a $4\times 4$ systolic array (2 rows shown) with PE row 0 having one outlier and three inliers in the $\mu$B and PE row 1 having all inlier weights. We assume $\mu$B=4, inlier and outlier quantization formats of MX-INT-2128 and MX-FP-44,4, respectively. To demonstrate that outlier partial sums processed through ReCoN results in accurate FP-outlier partial sum, we first show a $\mu B$ with one outlier and its actual FP-value post sharing of $\mu X$, having a value of $1.5_{10}$. If there were no outlier distribution and conditioning performed, the actual partial sum from an “outlier-specific PE” for the shown iAct=$32_{10}$ and iAcc=$8_{10}$ will be $56_{10}$. However as described in section 4.3, we condition and distribute outliers within a $\mu$B into Upper and Lower halves. Note: The hidden-bit is implicit and is only shown for demonstration. The $\mu$B after conditioning and distributing the outlier is mapped to PE row 0. Since PE row 1 does not have any outliers, it is directly mapped with its $\mu$B. PE row 0 computes partial sum outputs for inliers in columns 1 and 2, while it outputs $O_{\texttt{Upper},\texttt{Res}}$ and $O_{\texttt{Lower},\texttt{Res}}$ for the other two columns that have outliers. The output of PE row 0 is routed through ReCoN. In ReCoN, the inliers are passed down at all stages, since they are already present in their respective columns. The stage 1 switch at column 3, executes a swap operation and redirects the lower half outlier output towards the upper half while simultaneously sending iAcc through the other port. This iAcc corresponds to the iAcc input to column 3 PE in row 0. This is functionally correct because, the lower half of outlier is distributed in this column by zeroing out/pruning the inlier at that location, hence the partial sum output of that PE to the corresponding PE in the next row should only be iAcc. Similarly in stage 2 of ReCoN, the switch that receives the $O_{\texttt{Lower},\texttt{Res}}$ from stage 1 again executes a swap operation. Finally in stage 3 the outlier output halves are merged together to realize $O^{0}_{Res}$, while all other switches execute pass. This merge operation results in a $O^{0}_{Res}$ that matches the expected output as shown in the figure. The partial sums from ReCoN are then sent to PE row 1 which then perform MAC operations and sends it to the subsequent row since there are no outlier weights. Similarly all rows execute operations until the final partial sum is calculated.
@@ 6. Experimental Evaluations
6.1. Experimental Setup [^7] In all experimental figures and tables, we denote each quantization configuration by its $bb$ to enhance readability and clear comparisons.
@@
Models and Datasets. We benchmark on various FMs including LLMs and VLMs, for a comprehensive evaluation of our method. For LLMs, we evaluate on OPT (Zhang et al., 2022), LLaMA2 (Touvron et al., 2023) LLaMA3 (Meta, 2024), Phi-3 (SLM) (Abdin et al., 2024) and Mixtral (MoE) (Jiang et al., 2024) model families and the VLMs evaluated are OpenFlamingo (Awadalla et al., 2023), VILA (Lin et al., 2024b) and LlaVa (Liu et al., 2024a). We use a calibration dataset consisting of 128 random samples from the PILE dataset (Gao et al., 2020), so as to not overfit to any particular downstream domain (Lin et al., 2024a). We compare different LLMs based on perplexity (PPL $\downarrow$) on the WikiText2 (Merity et al., 2016) dataset and benchmark accuracy on 6 different tasks BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-c (Clark et al., 2018), MMLU (Hendrycks et al., 2020) and Winogrande (Sakaguchi et al., 2021). Similarly, we evaluate VLMs on 5 vision-language benchmarks: COCO captioning (Chen et al., 2015), VQAv2 (Goyal et al., 2017), VizWiz (Gurari et al., 2018), TextVQA (Singh et al., 2019), and GQA (Hudson and Manning, 2019).
Figure 8. Qualitative results of OpenFlamingo-9B (Awadalla et al., 2023) on 8-shot COCO captioning with weight-only quantization using OliVe (W4) and MicroScopiQ (W2). Text in red and green indicate incorrect and accurate captioning respectively.
@@
Algorithm Implementation. We implement the MicroScopiQ Quantization Framework in PyTorch (Paszke et al., 2019). All FMs are quantized using a single NVIDIA H100 GPU. The complete quantization process’ runtime ranges between 30 mins. to 9 hours depending on model size (3.8B to 175B), which is at par with recent SoTA quantization frameworks (Shao et al., 2023; Frantar et al., 2022).
@@
Algorithm Baselines. We compare the performance of our MicroScopiQ Quantization Framework against existing SoTA quantization algorithms: OmniQuant (Shao et al., 2023), AWQ (Lin et al., 2024a), GPTQ (Frantar et al., 2022) and co-design techniques: OliVe (Guo et al., 2023) and GOBO (Zadeh et al., 2020).
Figure 9. Weight-only quantization results for VLMs, OpenFlamingo-9B (Awadalla et al., 2023) on COCO and VILA-7B (Lin et al., 2024b) on VizWiz and TextVQA.
@@
Accelerator Implementation. We implement the MicroScopiQ accelerator in Verilog and synthesize, place and route using Synopsys Design Compiler, Innovus with a TSMC 7nm technology library to estimate area, power and latency (See accelerator floorplan in Figure 12(b)). All reported variants of MicroScopiQ at all scales attains a peak clock frequency of 1 GHz. We use CACTI (Muralimanohar et al., 2009) to estimate the area and power of on-chip memories. For end-to-end performance evaluation, we design a cycle accurate simulator based on DnnWeaver (Sharma et al., 2016) and BitFusion (Sharma et al., 2018), following previous works (Guo et al., 2023; Ramachandran et al., 2024b).
@@
Accelerator Baselines. We compare the MicroScopiQ accelerator with OliVe (Guo et al., 2023), ANT (Guo et al., 2022), GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) and AdaptivFloat (Tambe et al., 2020) across area, power and performance metrics. We also employ two NoC-based ML accelerators, MTIA (Firoozshahian et al., 2023) and Eyeriss-v2 (Chen et al., 2019), as baselines to demonstrate the minimal overhead of ReCoN on real-world industrial and academic accelerators respectively. All designs have a peak clock frequency of 1GHz. DeepScale (Sarangi and Baas, 2021) is employed to scale all designs to the 7nm process.
@@ 6.2. LLM Quantization Results
Weight-Only Quantization. In Table 4, we compare the WikiText2 PPL (lower the better) of different LLMs for W4A16 and W2A16 settings. Apparent from the table, MicroScopiQ consistently outperforms all the baselines across different models and quantization settings. At W4A16, MicroScopiQ achieves near-lossless quantization performance. MicroScopiQ has minimal increase in PPL due to its ability to quantize outliers at higher-precision and in FP-format to reduce outlier quantization error. Remarkably, MicroScopiQ outperforms/closely matches the SoTA training-based quantization framework OmniQuant across all models which requires 56×5-6$\times$5 - 6 × higher runtime than MicroScopiQ. At W2A16, group B techniques like AWQ and OliVe fail ($\geq 1e4$ PPL).
[…]
Benchmark Accuracy. In Figure 2(b), we compare the zero-shot task accuracy of W2A16 MicroScopiQ against W416 OliVe for three different LLM benchmarks. MicroScopiQ at lower-precision than OliVe achieves a remarkable $\geq$8higher accuracy on the benchmarks compared to OliVe. OliVe shows poor accuracy due to its flawed assumption on outlier locality, resulting in unintended outlier pruning. Similarly, in Table 4 we also compare MicroScopiQ against OliVe and OmniQuant at W2A16 setting on 4 other LLM benchmarks. MicroScopiQ consistently outperforms the algorithm SoTA OmniQuant by $3-4\%$ across all benchmarks.
Effective Bit Width. For weight $bb=2,4$, MicroScopiQ has an EBW of $4.15$ and $2.36$ bits respectively. Techniques like GOBO, has a very high EBW of $18.17$ and $15.6$ bits respectively. All other baselines employ software-managed metadata and have an EBW=$bb$.

### v3

@@ 5.3. Multi-Precision PE Array
Accumulation Stage. It receives the MUL-stage output (Res) and also iAcc from either ReCoN or the previous PE row. The controller signals the ADD stage via Outlier_Present to modify accumulation behavior based on whether inlier/outlier weights are present in the PE. If the Res corresponds to a multiplication of inlier weight with iAct, then the corresponding Res is added with iAcc using two adders with a multiplexer that enables multi-precision by propagating the carry from one adder to other. In low precision MODE, the two adders work independently to calculate partial sums in parallel. For high precision MODE, both adders work together to produce a single accumulation result with carry propagation through the multiplexer. On the other hand, if Res corresponds to multiplication of either of the outlier halves with iAct, the actual outlier accumulation is offloaded to ReCoN by concatenating Res and iAcc444the concatenated output is notated as $O_{\texttt{Upper},\texttt{Res}}$ or $O_{\texttt{Lower},\texttt{Res}}$.. This is done to prevent incorrect outlier partial sum calculation. The controller directs partial sum outputs to ReCoN or next PE row based on presence of outliers, via the OAcc_NoC/PE signal.
@@ 5.4. Redistribution and Coordination NoC
The Redistribution and Coordination NoC (ReCoN) is a multistage butterfly NoC in the MicroScopiQ accelerator time-multiplexed and shared across PE rows (see section 6.6). This is a more cost effective way of handling outliers compared to OliVe and GOBO (Guo et al., 2023; Zadeh et al., 2020) that handle outliers within the PE. Since PEs are large in number OliVe, GOBO will incur larger costs compared to MicroScopiQ.
@@
ReCoN Input Interface. ReCoN is shared and time-multiplexed across all PE rows, and at a particular instant many rows might require access to it. At the inputs, the column-wise arbiters guided by control signals from the controller, coordinate and resolve contention between different PE rows and ensure a fair scheduling (Note for simplicity in Figure 4, we conceptually depict this operation as a N-input multiplexer). Due to the skewed data flow of systolic array computations (Jouppi et al., 2017; Ramachandran et al., 2024b), partial sums from different columns of the same row do not arrive at the ReCoN input at the same cycle. A synchronization buffer (Figure 6(c)) is positioned between ReCoN and column-wise multiplexers which propagates the partial sums through sets of buffers of different lengths (column producing the partial sum fastest has the most buffers), synchronizing the arrival of inputs from different columns. The synchronization buffer sends an ACK signal to the PE row whose values have been accepted so that it can progress in its computation and the other simultaneously contesting rows will be on hold and acknowledged in the next cycle. When there are N simultaneously contesting rows, the row that is acknowledged the last adds an N-1 cycle latency to its processing. As we show in Figure 10(b), Figure 13(a) the num. of conflicts in accessing ReCoN are minimal and if latency is a critical factor multiple ReCoN units can be used to minimize conflicts.
ReCoN switch. It is a {2-input, 2-output} switch (Figure 6(c)) that performs three major operations based on a 3-bit configuration:
@@
Figure 7. End-to-end example illustrating the working of MicroScopiQ with a $4\times 4$ PE array (only two rows shown).
@@ 5.5. Post Processing and Output Scale Factor Computation
The shared output scale factor for $B_{M}$=128 is calculated as ${oAct_{sf}}={O_{sf}+iAct_{sf}}$ (See expansion of $O_{sf}$ in section 4.2.). The calculation of $oAct_{sf}$ (with simple adders and subtractors) is independent of the processing done in the PE array. Therefore, we overlap the computation of $oAct_{sf}$ with the processing of oActs to efficiently hide the computation latency similar to prior works (Dai et al., 2021; Ramachandran et al., 2024b). Since we maintain different scales for inliers and outliers (see section 4.2), and the oAct scale factor is calculated based on the outlier scale factor ($O_{sf}$) (we found this to result in least quantization error of oActs), the oActs which are generated through computation with only inlier weights (identified through the in/out control signal) are shifted by the shift value depicted in the “Post Processing and Scale Compute Unit” in Figure 4 to ensure conformity with the final scale factor. Finally, the oActs are scaled with the computed output scale factor through a simple right shift (since it is a power-of-two scale factor, division can be implemented through right shift operation) and quantized to MX-INT-(4/8)128 before being sent to external memory or routed back to the iAct buffer for computation with the next layer’s weights. Similar to prior works (Ramachandran et al., 2024b), (Keller et al., 2023) the post processing unit is also responsible for handling all non-linear operations.
@@
SmoothQuant (Xiao et al., 2023) 4/4 19.54 17.62 20.47 15.63 17.62 29.54 19.32 37.54 18.11 15.39
Atom (Zhao et al., 2024) 4/4 11.15 9.02 6.16 6.12 5.20 8.12 4.69 5.35 7.59 5.95
MicroScopiQ (Ours) 4/4 10.97 8.95 6.11 5.57 4.48 8.12 4.65 5.03 6.95 5.41
@@
Atom (Zhao et al., 2024) 2/8 11.95 10.13 9.23 8.54 6.33 9.13 6.35 6.14 7.46 7.29
MicroScopiQ (Ours) 2/8 11.77 9.98 9.06 8.06 6.33 9.08 6.02 6.17 7.38 6.82
Table 3. LLM quantization performance of MicroScopiQ augmented with OmniQuant (Omni-MicroScopiQ).
@@
Omni-MicroScopiQ 2/8 7.12 5.74 7.21
Table 4. Comparison of quantization performance for LLaMA2-70B on LLM benchmarks.
@@
MicroScopiQ 2/16 53.30 81.60 63.70 77.80
5.6. MicroScopiQ Walkthrough Example
In this section, we discuss an end-to-end example that concretely showcases the working of MicroScopiQ. In the example shown in Figure 7, we employ a $4\times 4$ systolic array (2 rows shown) with PE row 0 having one outlier and three inliers in the $\mu$B and PE row 1 having all inlier weights. Note: Not all rows require access to ReCoN, only rows that have one or more outlier weights (row 0) in the PE require access. We assume $\mu$B=4, inlier and outlier quantization formats of MX-INT-2128 and MX-FP-44,4, respectively.
1
To demonstrate that outlier partial sums processed through ReCoN results in accurate FP-outlier partial sum, we first show a $\mu B$ with one outlier and its actual FP-value after sharing of $\mu X$, and having a value of $1.5_{10}$. If there were no outlier distribution, the actual partial sum from an “outlier-specific PE” for the shown iAct=$32_{10}$ and iAcc=$8_{10}$ will be $56_{10}$. However as described in section 4.3, we distribute outliers within a $\mu$B into Upper and Lower halves. Note: The hidden-bit is implicit and is only shown for demonstration.
2
The $\mu$B after distributing the outlier is mapped to PE row 0. 3 PE row 0 computes partial sum outputs for inliers in columns 1 and 2, while it outputs $O_{\texttt{Upper},\texttt{Res}}$ and $O_{\texttt{Lower},\texttt{Res}}$ for the other two columns that have outliers. The output of PE row 0 is routed through ReCoN. In ReCoN, the inliers are passed down at all stages, since they are already present in their respective columns.
4
The stage 1 switch at column 3, executes a swap operation and redirects the lower half outlier output towards the upper half while simultaneously sending iAcc through the other port. This iAcc corresponds to the iAcc input to column 3 PE in row 0. This is functionally correct because, the lower half of outlier is distributed in this column by zeroing out/pruning the inlier at that location, hence the partial sum output of that PE to the corresponding PE in the next row should only be iAcc since the original weight at that location is zero. Similarly in stage 2 of ReCoN, the switch that receives the $O_{\texttt{Lower},\texttt{Res}}$ from stage 1 again executes a swap operation. Finally in stage 3 the outlier output halves are merged together to realize $O^{0}_{Res}$, while all other switches execute pass. This merge operation results in a $O^{0}_{Res}$ that matches the expected output.
5
The partial sums from ReCoN are then sent to PE row 1 which then perform MAC operations and sends it to the subsequent row since there are no outlier weights. Similarly all rows execute operations until the final partial sum is calculated.
@@ 6. Experimental Evaluations
6.1. Experimental Setup [^5] We denote each quantization configuration by its $bb$ to enhance readability and facilitate clear comparisons.
@@
Models and Datasets. We evaluate on OPT (Zhang et al., 2022), LLaMA2 (Touvron et al., 2023) LLaMA3 (Meta, 2024), Phi-3 (SLM) (Abdin et al., 2024) and Mixtral (MoE) (Jiang et al., 2024) LLM model families. The VLMs evaluated are OpenFlamingo (Awadalla et al., 2023), VILA (Lin et al., 2024b) and LlaVa (Liu et al., 2024a). We use a calibration dataset consisting of 256 random samples from the PILE dataset (Gao et al., 2020), so as to not overfit to any particular downstream domain (Lin et al., 2024a). We compare different LLMs based on perplexity (PPL $\downarrow$) on the WikiText2 (Merity et al., 2016) dataset and benchmark accuracy on 6 different tasks BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-c (Clark et al., 2018), MMLU (Hendrycks et al., 2020) and Winogrande (Sakaguchi et al., 2021). Similarly, we evaluate VLMs on 5 vision-language benchmarks: COCO captioning (Chen et al., 2015), VQAv2 (Goyal et al., 2017), VizWiz (Gurari et al., 2018), TextVQA (Singh et al., 2019), and GQA (Hudson and Manning, 2019).
@@
Algorithm Implementation. We implement the MicroScopiQ Quantization Framework in PyTorch (Paszke et al., 2019). All FMs are quantized using a single NVIDIA H100 GPU. The complete quantization process’ runtime ranges between 30 mins.–9 hours depending on model size (3.8B to 175B), which is at par with recent SoTA techniques (Frantar et al., 2022; Zhao et al., 2024).
@@
Algorithm Baselines. We compare the performance of our MicroScopiQ quantization framework against existing SoTA quantization algorithms: OmniQuant (Shao et al., 2023), AWQ (Lin et al., 2024a), GPTQ (Frantar et al., 2022), Atom (Zhao et al., 2024), SDQ (Jeong et al., 2024) and co-design techniques: OliVe (Guo et al., 2023), GOBO (Zadeh et al., 2020).
@@
Accelerator Implementation. The MicroScopiQ accelerator is implemented in Verilog RTL. We perform synthesis and PnR using Synopsys Design Compiler and Innovus with a TSMC 7nm technology library (See floorplan in Figure 11(b)) for area, power and latency calculations. All MicroscopiQ design variants attains a peak clock frequency of 1 GHz. We use CACTI (Muralimanohar et al., 2009) to estimate the area and power of on-chip memories. For end-to-end performance evaluation, we design a cycle accurate simulator based on DnnWeaver (Sharma et al., 2016) and BitFusion (Sharma et al., 2018), following previous works (Guo et al., 2023; Ramachandran et al., 2024b).
@@
Accelerator Baselines. We compare the MicroScopiQ accelerator against OliVe (Guo et al., 2023), ANT (Guo et al., 2022), GOBO (Zadeh et al., 2020), OLAccel (Park et al., 2018) and AdaptivFloat (Tambe et al., 2020) across area, power and performance metrics. To demonstrate the minimal overhead of ReCoN on NoC-based real-world industrial and academic accelerators, we also implement and compare against baseline MTIA (Firoozshahian et al., 2023) and Eyeriss-v2 (Chen et al., 2019). For a fair comparison, we ensure all designs attain a clock frequency of 1GHz. DeepScale (Sarangi and Baas, 2021) is employed to scale all designs to the 7nm process.
@@ 6.2. LLM Quantization Results
Weight-Only Quantization. In Table 4, we compare the WikiText2 PPL of different LLMs for W4A16 and W2A16 settings. MicroScopiQ consistently outperforms all the baselines across different models and quantization settings. At W4A16, MicroScopiQ achieves near-lossless quantization performance. Particularly at W2A16, the benefits of the MicroScopiQ method is evident, achieving up to a 2.04 decrease in PPL compared to the baselines. This is due to MicroScopiQ’s ability to quantize outliers at higher-precision and in FP-format to reduce quantization error. At W2A16, techniques like AWQ and OliVe fail ($\geq 1e4$ PPL, not depicted in table). MicroScopiQ outperforms the SoTA training-based framework OmniQuant across all models which requires 56×5-6$\times$5 - 6 × higher runtime.
Weight-Activation Quantization. Similarly, we compare the quantization performance of MicroScopiQ for W4A4 and W2A8 configurations in Table 4. Due to the dynamic nature of activations, very rarely do techniques quantize outliers in activation directly (Zhao et al., 2024). Instead, most techniques (Xiao et al., 2023; Shao et al., 2023) migrate the activation outliers to weights to enable simpler activation quantization. We borrow the migration strength ($\alpha$) hyper-parameter introduced in SmoothQuant (Xiao et al., 2023) to migrate the difficulty of quantizing outliers in activations to weights. We find that prior techniques (Xiao et al., 2023; Shao et al., 2023) can only migrate up to 50% of activation quantization difficulty to weights, i.e., $\alpha=0.5$ before beginning to introduce higher weight quantization error. MicroScopiQ’s robustness to higher presence of outliers in weights by effectively quantizing outliers at higher precision and identifying least important weights to prune, allows $\alpha$ as high as 0.7 i.e., migrating most of the activation outliers to weights. This enables simpler activation quantization with MX-INT-(4/8)128 and absolving the need to handle outliers in activations. With $\alpha=0.7$ for MicroScopiQ, $\alpha=0.5$ for SmoothQuant and learned migration strength for OmniQuant in Table 4, MicroScopiQ outperforms all baselines with up to 7.4$\times$ lower PPL across quantization settings. Furthermore, MicroScopiQ outperforms a recent work Atom (Zhao et al., 2024) that performs activation quantization with up to a 0.33 lower PPL.
Figure 8. Weight-only quantization for VLMs, OpenFlamingo-9B (Awadalla et al., 2023) on COCO and VILA-7B (Lin et al., 2024b) on VizWiz and TextVQA.
Benchmark Accuracy. In Figure 2(b), we compare the zero-shot task accuracy of W2A16 MicroScopiQ against W416 OliVe for three different LLM benchmarks. MicroScopiQ at lower-precision than OliVe achieves $\geq$8higher accuracy on the benchmarks compared to OliVe. OliVe shows poor accuracy due to its assumption on outlier locality, resulting in unintended outlier pruning. In Table 4 we also compare MicroScopiQ against OliVe and OmniQuant at W2A16 setting on 4 other LLM benchmarks. MicroScopiQ consistently outperforms all baselines by up to 9% across all benchmarks.
Effective Bit Width. For $bb=2,4$, MicroScopiQ has an EBW of $2.36$ and $4.15$ bits respectively. Techniques like GOBO, have a very high EBW of $15.6$, $18.17$ bits. All other baselines employ software-managed metadata and have an EBW=$bb$.